I am trying to figure out the length of a rectangular playing field that is 5 ft less than twice its width and the perimeter of the playing field is 230 ft, what is the length and width of the field?
Answer by checkley71(8403)

2(2X-5)+2X=230 OR 4X-10+2X=230 OR 6X=230+10 OR 6X=240 OR X=240/6 OR X=40 WHICH IS THE WIDTH AND 2*40-5=80-5=75
PROOF 2*40+2*75=230 OR 80+150=230 OR 230=230